by TEI TEI

08/03/16 – TTC was on site with John Gilbert yesterday troubleshooting the BMV area. It looks like its fed from AH-1 not AH-3 as originally thought. County to investigate and get back to TEI. Airflow on CV box was also locked in at 20% and several boxes were set to 0%, I would assume due to overcooling.

09/25/17 – No movement on this project. TTC sent an email to Tony about this project and helping if needed.